OpenStudy (anonymous):

What is the value of the variable that makes the statement true? a + (–7) + 3 = 0 A. –4 B. –3 C. 4 D. 7

OpenStudy (kendricklamar2014):

Step 1. Simplify both sides of the equation: a−7+3=0 = a−4=0 Step 2.Add 4 to both sides: a−4+4=0+4 a = 4 Answer: A = 4